OpenStudy (anonymous):

jack usually mows his lawn in 7 hours. Marilyn can mow the same yard in 3 hours. how much time would it take for them to mow the lawn together?

hero (hero):

If I remember correctly you do \[\frac{M \times N}{M+N}\]

hero (hero):

so 21/10 = 2.1 hrs